# Deep Semantic Inference over the Air: An Efficient Task-Oriented Communication System

### v1

@@ V Conclusion and Outlook
This work demonstrates that task-oriented communication, when implemented through deep learning model partitioning and semantic compression, can achieve a highly efficient balance between classification accuracy, computational load, and communication cost. When the model split point and semantic feature dimension are carefully aligned with device capabilities and channel conditions, resource usage can be significantly reduced while maintaining over 85% of baseline accuracy. The representative ResNet-18 and ResNet-34 models are adopted to conduct a systematic study on CIFAR-10 and CIFAR-100 datasets. This choice supports both reproducibility and generalizability across a wide range of classification tasks and deployment scenarios. The experimental results and analysis provide practical design insights for real-time inference systems in bandwidth- and energy-constrained environments.
This work suggests several promising directions for future research, including the development of adaptive task-oriented systems capable of dynamically adjusting model partitions and semantic feature dimensions in response to changing network conditions. Additionally, we aim to extend this framework to support more complex tasks, such as object detection and multi-modal reasoning, and to incorporate energy-aware and privacy-preserving mechanisms for real-world deployments.

### v2

@@ V Conclusion and Outlook
This work demonstrates that under the AWGN channel and varying dimensionalities, task-oriented communication implemented through deep learning model partitioning and semantic compression can achieve a highly efficient balance between classification accuracy, computational load, and communication cost. Experimental results with elaborative analysis provide practical design insights for real-time inference systems in bandwidth- and energy-constrained environments. Future research will focus on developing advanced feature compression and evaluation that dynamically adjusts model partitions and semantic feature dimensions in more realistic wireless scenarios. We also plan to extend this framework to more complex tasks, such as object detection and multi-modal reasoning, and to incorporate energy-aware and privacy-preserving mechanisms for real-world deployment.